OpenStudy (anonymous):

what is the surface area of a circle that has a diameter of 4 inches?

jhonyy9 (jhonyy9):

do you know formula for area of a circle ?

OpenStudy (anonymous):

nope i do not

OpenStudy (anonymous):

\[surface~ area~ of~ sphere=4\pi r^2\] \[Area~ of~ circle=\pi r^2\]

OpenStudy (anonymous):

\[radius~ r=\frac{ diameter }{2 }\]